# Function Induction and Task Generalization: An Interpretability Study with Off-by-One Addition

### v1

@@ Abstract
1 Introduction
As the capabilities of language models (LMs) continue to grow, users apply them to increasingly challenging and diverse tasks, accompanied by evolving expectations  [53, 36, 17]. Consequently, it becomes impractical to include every task of interest in a model’s training prior to deployment. In this context, task-level generalization—the ability of a model to perform novel tasks at inference time—becomes highly crucial and valued.
Prior work shows that LMs already exhibit this capability to a significant extent through in-context learning [3, 5, 24]. The underlying mechanisms of this behavior are being actively investigated, with work on induction heads [30] and function vectors [12, 37] offering substantial insights. However, our understanding is still limited, especially regarding more complex generalization scenarios involving unexpected elements or newly defined concepts in the task.
In this work, we aim to enhance our understanding of how models handle novelty and unconventionality with one counterfactual task: off-by-one addition (i.e., 1+1=3, 2+2=5, 3+3=?). For humans, this task consists of two sequential steps: standard addition, followed by an unexpected increment of one to the sum. When a language model is prompted to perform this task with in-context learning, we anticipate two possible outcomes: (1) the model acquires the intended +1 operation and thus outputs 7, or (2) it adheres to fundamental arithmetic principles and outputs 6.
We begin our study by evaluating six contemporary LMs on off-by-one addition. Our findings indicate that all evaluated models consistently demonstrate the first outcome, effectively leveraging in-context examples; furthermore, performance increases consistently as more shots are used. Motivated by these observations, we seek a more comprehensive understanding of how models perform off-by-one addition, and in particular, the +1 step of the task. To this end, we employ mechanistic interpretability and path patching techniques [42], which enables us to trace the model’s output logits to a specific set of attention heads and their interconnections responsible for +1 behavior.
Our analysis with Gemma-2 (9B) [8] reveals that the model’s computation of +1 is mainly governed by three groups of attention heads. Notably, two of these groups and their connections resemble the structure of the induction head mechanism described in prior work [30] [^2] Induction heads facilitate a language model’s pattern copying behavior in sequences like [A][B]...[A] $\rightarrow$ [B]. See Appendix A for further details.. This observation leads to our hypothesis of a function induction mechanism—a generalization of the induction head mechanism that operates at the function level. Our analysis also reveals that the +1 function is transmitted along six (or more) paths in the model’s computation graph; in each path, an attention head writes a distinct fraction of the function, whose aggregate effect yields the complete +1 function.
We further validate the universality of our findings across models and tasks [29, 23]. Regarding models, we repeat the path patching procedure with Mistral-v0.1 (7B) [16] and Llama-3 (8B) [38], confirming the existence of the function induction mechanism, though in slightly varied forms. Regarding tasks, we extend our analysis with four task pairs—off-by-$k$ addition, shifted multiple-choice QA, Caesar Cipher, and base-8 addition—designed to replace sub-steps in off-by-one addition with substantially different operations. We demonstrate the reuse of the same mechanism in these task pairs. Overall, our results highlight the flexible and composable nature of the function induction mechanism we have characterized, and provide an improved understanding of how language models may generalize when encountering unexpected aspects in a task.
@@ 2 LMs Learn Off-by-One Addition in Context
Off-by-one addition is a synthetic, counterfactual task involving two steps. The first step is standard addition, and the second, unexpected step is a +1 function. In our work, we are interested in whether and how the model can perform this task with in-context learning. We provide concrete 4-shot examples of standard addition and off-by-one addition in Table 1. In this section, we first evaluate contemporary language models on this tasks and describe our observations.
@@
Data.
To create the evaluation data, we randomly sample 100 test cases, each with 32 in-context examples ($a_{i}+b_{i}=c_{i}$) and one test example ($a_{test}+b_{test}=c_{test}$). We sample $a,b,c$ from the range of [0,999], and restrict that for all $i$, $c_{test}\neq c_{i}$. This is to make sure these test cases evaluate models on inducing +1 function, instead of copying and pasting the answer ($c_{test}$) from the previous context ($c_{i}$).
Models.
We evaluate six recent LMs on this task: Llama-2 (7B) [38], Mistral-v0.1 (7B) [16], Gemma-2 (9B) [8], Qwen-2.5 (7B) [48], Llama-3 (8B) [9] and Phi-4 (14B) [1]. These models were developed by different organizations, employ different number tokenization methods, and were released in different years, thereby providing a diverse and representative sample. Please refer to Table 4 for details of these models.
Evaluation Results.
In Fig. 1, we report the accuracy when different numbers of in-context examples are used. All evaluated models exhibit non-trivial performance on this task, demonstrating that this behavior is pervasive. Additionally, performance always improves as the number of shots increases, indicating effective utilization of the in-context examples. Notably, more recent models like Llama-3 (8B) and Phi-4 (14B) achieve the strongest performance, with near perfect results in the 8-shot experiments. More details of our evaluation (e.g., reporting accuracy on standard addition, using a smaller number range like [0,9], or removing the restriction of $c_{test}\neq c_{i}$) are deferred to §B.
@@ 3.1 Background: Mechanistic Interpretability and Path Patching
Mechanistic interpretability is a subfield of interpretability that aims to reverse-engineer model computations and establish “correspondence between model computation and human-understandable concepts.” [42] A transformer-based language model can be viewed as a computation graph $M$, where components like attention heads and MLP layers serve as nodes, and their connections as edges. We use $M(y|x)$ to denote the logit of token $y$ when using $x$ as the input prompt. A circuit $C$ is a subgraph of $M$ that is responsible for a certain behavior. In our study, the behavior of interest is the induction and application of the +1 function in off-by-one addition.
The specific method we rely on is path patching [42], which generalizes activation patching [22] and causal mediation [40] from prior work. In the past, such technique has supported interpretability findings on a wide range of model behaviors [11, 35, 33, 19].
@@
To simplify the notation, we define $F(C,x)$ as the logit difference between $y_{base}$ (6) and $y_{cont}$ (7) when prompted with $x$ and using the circuit $C$ while knocking out nodes outside $C$ in the computation graph, i.e., $F(C,x)=C(y_{base}|x)-C(y_{cont}|x)$. Following Wang et al. [42], we quantify the effect of a replacement by first computing $F(M^{\prime},x_{cont})$, and then normalize it by the logit difference before intervention, i.e., $r=\frac{F(M^{\prime},x_{cont})-F(M,x_{cont})}{F(M,x_{cont})-F(M,x_{base})}$. See §C.1 for its expansion and explanations. The resulting ratio $r$, which we refer to as relative logit difference, will typically fall in the range of [-100%, 0%], with -100% representing the model favors $y_{base}$ (i.e., the model losts its ability on off-by-one addition after replacement), and 0% representing the model favors $y_{cont}$.
@@ 3.2 Circuit Discovery
Refer to captionRefer to caption
@@
(d) Group 2 heads attend to the answer token ($c_{i}$) of previous ICL examples at the position of “=”.
@@
Patching to the Output Logits.
Our investigation begins by setting the output logits as the target node, effectively asking “which attention heads directly influence the model output?” The results, visualized in Fig. 2(a), highlight 10 attention heads with a relative logit difference $|r|>2\%$.
We further investigate the attention pattern of the highlighted heads and categorized them into two groups. Group 1 heads appear exclusively in the last two layers of the model, and mainly attend to the current token and the <bos> token at each position (Fig. 2(c)). Group 2 heads present periodical patterns consistent with the ICL examples in the prompt (Fig. 2(d)). Specifically, at the position of the last “=” token, where the model is expected to generate the answer as the next token, these attention heads will attend to the answer token ($c_{i}$) in previous ICL examples ($a_{i}+b_{i}=c_{i}$).
@@
Patching to the Value of Group 2 Heads.
To further trace down the origin of the +1 function, we set the value of each head in Group 2 as the target node for path patching. For example, H39.7 (Head 7 in Layer 39) is a representative head in Group 2 with a relative logit difference $r$ of $-27\%$ when patching to the final output. When setting H39.7’s value as the target node and performing path patching, three heads are highlighted (Fig. 2(b)) and all of these heads follow the pattern of attending to the previous token at certain positions (Fig. 2(e)). In particular, these head attend to the “=” token from the answer token $c_{i}$ in each in-context example. We repeat this procedure for remaining heads in Group 2 and identify more attention heads with the previous-token attending behavior. We collectively refer to them as Group 3 heads.
@@
The Function Induction Hypothesis.
Fig. 3 provides an overview of the circuit we identified, illustrating the connections of the three head groups and highlighting the token positions they operate on. The comprehensive list of heads in each group can be found in §C.2.1 and Fig. 4(c).
We find it particularly intriguing that the structure of the circuit, in particular Group 2 and Group 3, resembles the structure of induction heads [30], a known mechanism responsible for language model’s copy-paste behavior. In the induction head mechanism, a previous token head “copies information from the previous token to the next token”, and an induction head “uses that information to find tokens preceded by the present token.” [30] We provide an illustration for the original induction head mechanism in §A. We hypothesize that the circuit we identify generalizes this known mechanism from the token-level to the function-level. Based on this intuition, the three groups of attention heads will cooperate as follows:
@@
4 Circuit Evaluation and Analysis
@@
Previously, we constructed the function induction hypothesis based on our path patching results and its structural similarity to that of the induction heads mechanism. In this section, we dive deeper into the properties of the identified circuit by evaluating its quality and conducting additional analyses, with the goal of providing a more granular understanding of its behavior.
Figure 4: Circuit Evaluation.
Circuit Evaluation.
@@
We begin our evaluation with head ablation, a common technique to validate a head’s involvement in a specific model behavior [10, 46]. Here, we focus on FI heads and “ablate” a head by replacing its output in the forward pass on $x_{cont}$ with the corresponding head output in the forward pass on $x_{base}$. As shown in Fig. 4(a), the complete, unablated model achieved an accuracy of 86% on 16-shot off-by-one addition. Upon ablating the six FI heads, the model’s behavior switched back to standard addition, resulting in 100% accuracy on standard addition and 0% on off-by-one addition. For a controlled comparison, we also ablated six randomly selected heads; these showed minimal influence on either the base or contrast accuracy. This set of results provides preliminary evidence that the six FI heads are necessary in off-by-one addition.
We further conduct more rigorous evaluation of our circuit on the faithfulness, completeness, and minimality criteria introduced in [42]. Note that we focus on interpreting the “off-by-one” component of the task, rather than the standard addition component. Hence, these circuit evaluation metrics are adapted accordingly. The faithfulness metric measures whether a circuit $C$ has a similar performance to the full model $M$, i.e., whether $F(C,x_{cont})$ is close to $F(M,x_{cont})$, with $F(C,x)$ defined earlier in §3.1. We find that $F(M,x_{base})=7.17$, $F(M,x_{cont})=-1.26$, and $F(C,x_{cont})=0.56$, suggesting that $C$ recovers $\frac{7.17-0.56}{7.17-(-1.26)}=78.4\%$ of the performance of $M$.
The completeness criterion evaluates whether for each subset $K\subseteq C$, the difference between $F(C\backslash K,x_{cont})$ and $F(M\backslash K,x_{cont})$ is small. In the following, we will omit the $x_{cont}$ term for brevity. We use various different sets (e.g., randomly or greedily selected) as $K$ and report the results in Fig. 4(b). We find most points representing $(F(C\backslash K),F(M\backslash K))$ fall slightly below the $x=y$ line, while maintaining a monotonic trend, suggesting that the circuit $C$ is partially complete. This represents the best we can achieve with our current methodology. We also find that when $K$ is the set of all PT heads or all FI heads, both $f(C\backslash K)$ and $f(M\backslash K)$ are high, suggesting that the model favors $y_{base}$ in next-token generation (i.e., 3+3=6) and switches back to standard addition under these ablation conditions. These observations are consistent with our function induction hypothesis.
Lastly, the minimality criterion measures whether each head $v$ in $C$ is necessary, by seeking a subset $K\subseteq C\backslash\{v\}$ that has a high score of $|F(C\backslash(K\cup\{v\}))-F(C\backslash K)|$. We manually constructed the $K$ sets for this purpose. As shown in Fig. 4(c), each head in $C$ is relevant to the task and has a non-trivial effect (>2%) in performing off-by-one addition.
@@
Figure 5: Individual and Overall Effect of Identified FI Heads. Each head writes out different information, which aggregates to implement the function of $f(x)=x+1$. Rescaled and clearer plots for heads H32.6, H25.13, and H32.4 are provided in §D.
What do FI heads write out?
@@
In our hypothesis, FI heads are responsible for writing the +1 function to the residual stream at the “=” token. This behavior is highly relevant to recent findings on function vectors in language models [37, 12], which indicates that a small number of attention heads effectively transport task representations (i.e., function vectors) in in-context learning. The FI heads we identified align with this description, and moreover, generalize these findings because the mechanism we discovered operates within a multi-step task.

### v2

@@ Abstract
Large language models demonstrate the intriguing ability to perform unseen tasks via in-context learning. However, it remains unclear what mechanisms inside the model drive such task-level generalization. In this work, we approach this question through the lens of off-by-one addition (i.e., 1+1=3, 2+2=5, 3+3=?), a two-step, counterfactual task with an unexpected +1 function as a second step. Leveraging circuit-style interpretability techniques such as path patching, we analyze the models’ internal computations behind their performance and present three key findings. First, we uncover a function induction mechanism that explains the model’s generalization from standard addition to off-by-one addition. This mechanism resembles the structure of the induction head mechanism found in prior work and elevates it to a higher level of abstraction. Second, we show that the induction of the +1 function is governed by multiple attention heads in parallel, each of which emits a distinct piece of the +1 function. Finally, we find that this function induction mechanism is reused in a broader range of tasks, including synthetic tasks such as shifted multiple-choice QA and algorithmic tasks such as base-8 addition. Overall, our findings offer deeper insights into how reusable and composable structures within language models enable task-level generalization. [^1] Code: INK-USC/function-induction
1 Introduction
As the capabilities of language models (LMs) continue to grow, users apply them to increasingly challenging and diverse tasks, accompanied by evolving expectations (Zhao et al., 2024; Tamkin et al., 2024; Kwa et al., 2025). Consequently, it becomes impractical to include every task of interest in a model’s training prior to deployment. In this context, task-level generalization—the ability of a model to perform novel tasks at inference time—becomes highly crucial and valued.
Prior work shows that LMs already exhibit this capability to a significant extent through in-context learning (Brown et al., 2020; Chen et al., 2022; Min et al., 2022a). The underlying mechanisms of this behavior are being actively investigated, with work on induction heads (Olsson et al., 2022) and function vectors (Hendel et al., 2023; Todd et al., 2024) offering substantial insights. However, our understanding is still limited, especially regarding more complex generalization scenarios involving unexpected elements or newly defined concepts in the task.
In this work, we aim to enhance our understanding of how models handle novelty and unconventionality with one counterfactual task: off-by-one addition (i.e., 1+1=3, 2+2=5, 3+3=?). For humans, this task consists of two sequential steps: standard addition, followed by an unexpected increment of one to the sum. When a language model is prompted to perform this task with in-context learning, we anticipate two possible outcomes: (1) the model acquires the intended +1 operation and thus outputs 7, or (2) it adheres to fundamental arithmetic rules and outputs 6.
We begin our study by evaluating six contemporary LMs on off-by-one addition. Our findings indicate that all evaluated models consistently demonstrate the first outcome, effectively leveraging in-context examples; furthermore, performance increases consistently as more shots are used. Motivated by these observations, we seek a more comprehensive understanding of how models perform off-by-one addition, and in particular, the +1 step of the task. To this end, we employ mechanistic interpretability and path patching techniques (Wang et al., 2023), which enables us to trace the model’s output logits to a specific set of attention heads and their interconnections responsible for +1 behavior.
@@ 1 Introduction
We further validate the universality of our findings across models and tasks (Olah et al., 2020; Merullo et al., 2024). Regarding models, we repeat our analysis on Mistral-v0.1 (7B) (Jiang et al., 2023), Llama-2 (7B) (Touvron et al., 2023) and Llama-3 (8B) (Touvron et al., 2023), confirming the existence of the function induction mechanism, though in slightly varied forms. Regarding tasks, we extend our analysis with four task pairs—off-by-$k$ addition, shifted multiple-choice QA, Caesar Cipher, and base-8 addition—designed to replace sub-steps in off-by-one addition with substantially different operations. We demonstrate the reuse of the same mechanism in these task pairs.
Overall, our results advance our understanding of important language model capabilities such as in-context learning and latent multi-step reasoning. They highlight the flexible and composable nature of the function induction mechanism we have characterized, and provide substantive insights into how models may generalize when encountering novel task variations.
@@ 2 LMs Learn Off-by-One Addition in Context
Off-by-one addition is a synthetic, counterfactual task involving two steps. The first step is standard addition, and the second, unexpected step is a +1 function. In this work, we are interested in whether and how the model can perform this task with in-context learning. We provide concrete 4-shot examples of standard addition and off-by-one addition in Table 1. In this section, we first evaluate contemporary language models on this task and describe our observations.
@@
Data. To create the evaluation data, we randomly sample 100 test cases, each with 32 in-context examples ($a_{i}+b_{i}=c_{i}$) and one test example ($a_{test}+b_{test}=c_{test}$). We sample $a,b,c$ from the range of [0,999], and restrict that for all $i$, $c_{test}\neq c_{i}$. This is to make sure these test cases evaluate models on inducing +1 function, instead of copying and pasting the answer token ($c_{test}$) from the previous context ($c_{i}$).
Models. We evaluate six recent LMs on this task: Llama-2 (7B) (Touvron et al., 2023), Mistral-v0.1 (7B) (Jiang et al., 2023), Gemma-2 (9B) (Gemma Team et al., 2024), Qwen-2.5 (7B) (Yang et al., 2024a), Llama-3 (8B) (Grattafiori et al., 2024) and Phi-4 (14B) (Abdin et al., 2024). These models were developed by different organizations, employ different number tokenization methods, and were released in different years, thereby providing a diverse and representative sample. Please refer to Table 4 for details of these models.
@@
Evaluation Results. In Fig. 1, we report the accuracy when different numbers of in-context examples are used. All evaluated models exhibit non-trivial performance on this task, demonstrating that this behavior is pervasive. Additionally, performance always improves as the number of shots increases, indicating effective utilization of the in-context examples. Notably, more recent models like Llama-3 (8B) and Phi-4 (14B) achieve the strongest performance, with near perfect results in the 8-shot experiments. More details of our evaluation (e.g., reporting accuracy on standard addition, using a smaller number range like [0,9], or removing the restriction of $c_{test}\neq c_{i}$) are deferred to §C.
@@ 3.1 Background: Mechanistic Interpretability and Path Patching
Mechanistic interpretability is a subfield of interpretability that aims to reverse-engineer model computations and establish “correspondence between model computation and human-understandable concepts.” (Wang et al., 2023) A transformer-based language model can be viewed as a computation graph $M$, where components like attention heads and MLP layers serve as nodes, and their connections as edges. We use $M(y|x)$ to denote the logit of token $y$ when using $x$ as the input prompt. A circuit $C$ is a subgraph of $M$ that is responsible for a certain behavior. In our study, the behavior of interest is the induction and application of the +1 function in off-by-one addition.
The specific method we rely on is path patching (Wang et al., 2023), which is built on activation patching (Meng et al., 2022) and causal mediation (Vig et al., 2020) methods from prior work. In the past, such technique has supported interpretability findings on a wide range of model behaviors (Hanna et al., 2023; Stolfo et al., 2023; Prakash et al., 2024b; Li et al., 2025).
@@
To simplify the notation, we define $F(C,x)$ as the logit difference between $y_{base}$ (6) and $y_{cont}$ (7) when prompted with $x$ and using the circuit $C$ while knocking out nodes outside $C$ in the computation graph, i.e., $F(C,x)=C(y_{base}|x)-C(y_{cont}|x)$. Following Wang et al. (2023), we quantify the effect of a replacement by first computing $F(M^{\prime},x_{cont})$, and then normalize it by the logit difference before intervention, i.e., $r=\frac{F(M^{\prime},x_{cont})-F(M,x_{cont})}{F(M,x_{cont})-F(M,x_{base})}$. See §D.1 for its expansion and explanations. The resulting ratio $r$, which we refer to as relative logit difference, will typically fall in the range of [-100%, 0%], with -100% representing the model favors $y_{base}$ (i.e., the model losts its ability on off-by-one addition after replacement), and 0% representing the model favors $y_{cont}$.
@@ 3.2 Circuit Discovery
Patching to the Output Logits. Our investigation begins by setting the output logits as the target node, effectively asking “which attention heads directly influence the model output?” The results, visualized in Fig. 2(a), highlight 10 attention heads with a relative logit difference $|r|>2\%$.
[Uncaptioned image][Uncaptioned image]
@@
(d) Group 2 heads attend to the answer tokens ($c_{i}$) of previous ICL examples at the position of “=”.
@@
We further investigate the attention pattern of the highlighted heads and categorized them into two groups. Group 1 heads appear exclusively in the last two layers of the model, and mainly attend to the current token and the <bos> token at each position (Fig. 2(c)). Group 2 heads present periodical patterns consistent with the ICL examples in the prompt (Fig. 2(d)). Specifically, at the position of the last “=” token, where the model is expected to generate the answer as the next token, these attention heads will attend to the answer tokens ($c_{i}$) in previous ICL examples ($a_{i}+b_{i}=c_{i}$).
@@
Patching to the Value of Group 2 Heads. To further trace down the origin of the +1 function, we set the value of each head in Group 2 as the target node for path patching. For example, H39.7 (Head 7 in Layer 39) is a representative head in Group 2 with a relative logit difference $r$ of $27\%$ when patching to the final output. When setting H39.7’s value as the target node and performing path patching, three heads are highlighted (Fig. 2(b)) and all of these heads follow the pattern of attending to the previous token at certain positions (Fig. 2(e)). In particular, at the answer token $c_{i}$ in each in-context example, these head attend to the “=” token immediately before $c_{i}$. We repeat this procedure for remaining heads in Group 2 and identify more attention heads with the previous-token attending behavior. We collectively refer to them as Group 3 heads.
@@
The Function Induction Hypothesis. Fig. 3 provides an overview of the circuit we identified, illustrating the connections of the three head groups and highlighting the token positions they operate on. The comprehensive list of heads in each group can be found in §D.2.1 and Fig. 21(b).
We find it particularly intriguing that the structure of the circuit, in particular Group 2 and Group 3, resembles the structure of induction heads (Olsson et al., 2022), a known mechanism responsible for language model’s copy-paste behavior. In the induction head mechanism, a previous token head “copies information from the previous token to the next token”, and an induction head “uses that information to find tokens preceded by the present token.” (Olsson et al., 2022) We provide an illustration for the original induction head mechanism in §B.
@@
We hypothesize that the circuit we identify generalizes this known mechanism from the token-level to the function-level. Based on this intuition, the three groups of attention heads will cooperate as follows:
@@
4 Circuit Validation and Analysis
@@
Previously, we constructed the function induction hypothesis based on our path patching results and its structural similarity to that of the induction heads mechanism. In this section, we dive deeper into the identified circuit, aiming to provide a more granular understanding.
Figure 4: Head Ablation Results.
@@
Initial Validation: Ablating FI Heads. We begin our investigation with head ablation, a common technique to validate a head’s involvement in a specific model behavior (Halawi et al., 2023; Wu et al., 2025). Here, we focus on FI heads and “ablate” a head by replacing its output in the forward pass on $x_{cont}$ with the corresponding head output in the forward pass on $x_{base}$. As shown in Fig. 4(a), the complete, unablated model achieved an accuracy of 86% on 16-shot off-by-one addition. Upon ablating the six FI heads, the model’s behavior switched back to standard addition, resulting in 100% accuracy on standard addition and 0% on off-by-one addition. For a controlled comparison, we also ablated six randomly selected heads; these showed minimal influence on either the base or contrast accuracy. This set of results provides preliminary evidence that the six FI heads are necessary in off-by-one addition.
@@
Figure 5: Individual and Overall Effect of Identified FI Heads. Each head writes out different information, which aggregates to implement the function of $f(x)=x+1$ (bottom-right panel). (*) Effects of H32.6, H25.13, and H32.4 are rescaled to [-0.15, 0.15] to make the patterns more readable.
@@
Further Validation: Measuring the Causal Effect of FI Heads. In our hypothesis, FI heads are responsible for writing the +1 function to the residual stream at the “=” token. This behavior is highly relevant to recent work (Todd et al., 2024; Hendel et al., 2023) which indicates that a small number of attention heads (i.e., function vector heads) effectively transport task representations (i.e., function vectors) in in-context learning. The FI heads we identified align with this description, and moreover, uncover a novel instantiation of the mechanism that operates within multi-step tasks.
The notion of function vectors inspires us to further validate the role of FI heads through their causal effect on a naive prompt $x_{naive}$, e.g., $2\n$3=?”, for which the model is expected to assign a high probability to “3”. If a FI head indeed writes out the +1 function, adding its output to the residual stream at the final “=” token should cause the model to increase its probability of generating “4”.
Concretely, we construct the naive prompt ${x-1}\n${x}=?” for $x\in[0,9]$, and track the model’s logits for tokens $[0,9]$ both before and after adding the FI head output to the residual stream at the corresponding layer. This leads to a $10\times 10$ heatmap, where the value at cell $(x_{input},y_{output})$ represents the change in logits for token $y$ when the function vector is added.
In Fig. 5, we present these heatmaps for each of the six FI heads identified in §3.2. We include three additional heads (H32.4, H28.6, H24.9) that, while showing modest effects ($1\%<|r|<2\%$) in §3.2, contribute meaningfully to the +1 function as revealed by this analysis. We find that FI heads work collaboratively—each of them contributes a distinct piece to the overall +1 function. For example, with an input $x$, H39.7 promotes $x+1$, H28.6 suppress x1x-1, H32.1 promotes digits greater than $x$, H24.9 suppresses $x$. When the outputs of these nine heads are added to the final residual stream altogether, their combined effect implements the +1 function, as depicted in the last panel of Fig. 5.
Universality of Function Induction. To investigate the universality of our findings across models, we repeat the path patching experiments with Llama-3 (8B), Llama-2 (7B), and Mistral-v0.1 (7B). We identified all three groups of heads across these models, except that the two consolidation heads identified in Mistral-v0.1 display weaker and less consistent signals. Still, these observations provide promising evidence that the function induction mechanism is general and consistently emerges across various language models. See §D.2 for more details.
FI Heads (Ours) and FV Heads (Todd et al., 2024) are Two Disjoint Sets of Heads. While our analysis demonstrates that FI heads transport task representations similarly to the FV heads described in prior work, a direct comparison with Llama 2 (7B) reveals important distinctions. Todd et al. (2024) reported that FV heads appear in early-middle layers of the model (before layer 20), whereas our FI heads are located in late layers of the model (layer 29-31). There is no overlap between the two sets of heads, suggesting that our work presents a distinct, previously undocumented finding. We hypothesize that FI heads can be seen as an instantiation of the broader FV head mechanism, but are only triggered in multi-step tasks where late layers are used to perform the late steps. See §D.2.4 for the full list of FI/FV heads and §6 for related discussions.
Additional Analysis. Due to space limits, we defer various supporting evidence to the appendix. We conduct a rigorous evaluation of our circuit using the faithfulness, completeness, and minimality criteria introduced in Wang et al. (2023). Our circuit mostly satisfies these criteria, and we discuss the results in §E. We deliberately focus on FI heads in §4 given the interesting insights from these results. We provide further validation and analysis of consolidation heads and previous token heads in §F.
5 Task Generalization with Function Induction
Our investigation so far suggests that function induction is the key mechanism enabling the model to generalize from standard addition and manage the unexpected +1 step in off-by-one addition. Given the importance of task generalization for capable AI systems, we aim to explore the broader usage of this mechanism. In this section, we investigate the role of function induction in a range of synthetic and algorithmic tasks. Specifically, §5.1 introduces the four task pairs examined, and §5.2 presents the overall findings and additional analyses for two of these pairs.
5.1 Tasks